The Rat River flows at a rate of 5 km/h. In order for a boat to travel 75 km upriver and then return in a total of 8 hr, how fast must the boal travel in still water?
Answer by checkley77(12844)

D=R*T
T=D/R
8=75/(R-5)+75/(R+5)
8=[75(R+5)+75(R-5)]/(R-5)(R+5)
8=[75R+375+75R-375]/(R^2-25)
8R^2-200=150R
8R^2-150R-200=0
(8R+10)(R-20)=0
R-20=0
R=20 KM/H IS THE SPEED OF THE BOAT.
PROOF:
8=75/15+75/25
8=5+3
8=8